OpenStudy (anonymous):

Simplify. (4x2 + 7x - 5) - (3x2 - 2x - 4) A. x2 + 9x - 1 B. -x2 + 5x + 1 C. -x2 - 9x - 9 D. 7x2 + 5x - 9

OpenStudy (anonymous):

The Answer Is A .

OpenStudy (hoblos):

distribute the (-) in the second expression (4x2 + 7x - 5) - (3x2 - 2x - 4)= 4x2 + 7x - 5 - 3x2 + 2x + 4 now combine like terms to get x2 + 9x - 1

OpenStudy (anonymous):

4x^2 - 3x^2 + 7x + 2x - 5 + 4 x^2 + 9x - 1 so A.

OpenStudy (anonymous):

4x^2-3x^2=x^2 7x-(-2x)= 7x+2x= 9 -5-(-4)=-5+4=-1 x^2+9-1